OpenStudy (anonymous):

A thief is spotted by a policeman from a distance of 100 m. When the policeman starts the chase, the thief also starts running. If the speed of thief be 4 kmh" 1 and that of the policeman 10 kmh 1 , how far will the thief have to run before he is caught?

OpenStudy (zpupster):

lets find the relative speed: police speed - thief speed = relative speed 10 - 4 = 6 km /hr= 6000m/hr figure out the time taken for 100 m 100m/6000m/ hr = 1/60 hr = 1 minute distance traveled by thief in one minute 4000m/hr)((1)/60hr)= 4000/60 = 66.67 m